use elimination to solve the system 4x+7y=-14 and 8x+5y=8 for x
Answer by checkley71(8403)

4X+7Y=-14 MULTIPY THIS EQUATION BY -5 WE GET -20X-35Y=70
8X+5Y=8 NOW MULTIPLY THIS EQUATION BY 7 WE GET 56X+35Y=56 NOW ADD
-20X-35Y=70
56X+35Y=56
-----------------------------------------
36X=126
X=126/36
X=3.5 SOLUTION